OpenStudy (anonymous):

find the value of sin^2 π/12, without using calculator. Thanks.

ganeshie8 (ganeshie8):

use the double angle formula

ganeshie8 (ganeshie8):

\(\cos 2\theta = 1-2\sin^2\theta\)

OpenStudy (anonymous):

got it! thanks!!!

ganeshie8 (ganeshie8):

np :)